rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#10892 — FS#14792 — bhs2-20a/b-n56

Attached to Project— Network
Incident
Beauharnois → BHS-2
Planned
0%
Both switches that are working in pair have rebooted due to an HA process dysfunction. We are working to have them up as fast as possible.
Comment by OVH - Thursday, 24 September 2015, 15:25PM

All fex are up now, traffic has to go back on all customer interfaces.